OpenStudy (anonymous):

Solve the following square roots when x>-3. (Below)

OpenStudy (anonymous):

\[\sqrt{(x-2)^{2}} \sqrt{(x+2)^2} \sqrt{(x+5)^2}\]

OpenStudy (anonymous):

\[ \sqrt{(x-2)^{2}} \sqrt{(x+2)^2} \sqrt{(x+5)^2}=\sqrt{(-3-2)^{2}} \sqrt{(-3+2)^2} \sqrt{(-3+5)^2}=\\ \sqrt{(-5)^{2}} \sqrt{(-1)^2} \sqrt{(2)^2}=5(1)(2)=10 \]

OpenStudy (anonymous):

@eliassaab Thank you! They were actually individual problems, but your steps definitely helped me!

OpenStudy (anonymous):

YW, you should put \\ between them to have each one on a line by itself.

OpenStudy (anonymous):

Sorry, I was having trouble with spacing them, thanks for the tip :)

OpenStudy (anonymous):

to put space between them us \quad